# Instructions for Authors: Title in Title Case

### v1

@@ Abstract
Traditional imitation learning focuses on modeling the behavioral mechanisms of experts, which requires a large amount of interaction history generated by some fixed expert. However, in many streaming applications, such as streaming recommender systems, online decision-makers typically engage in online learning during the decision-making process, meaning that the interaction history generated by online decision-makers includes their behavioral evolution from novice expert to experienced expert. This poses a new challenge for existing imitation learning approaches that can only utilize data from experienced experts. To address this issue, this paper proposes an inverse batched contextual bandit (IBCB) framework that can efficiently perform estimations of environment reward parameters and learned policy based on the expert’s behavioral evolution history. Specifically, IBCB formulates the inverse problem into a simple quadratic programming problem by utilizing the behavioral evolution history of the batched contextual bandit with inaccessible rewards. We demonstrate that IBCB is a unified framework for both deterministic and randomized bandit policies. The experimental results indicate that IBCB outperforms several existing imitation learning algorithms on synthetic and real-world data and significantly reduces running time. Additionally, empirical analyses reveal that IBCB exhibits better out-of-distribution generalization and is highly effective in learning the bandit policy from the interaction history of novice experts.
@@ 1 Introduction
Traditional imitation learning (IL) focuses on learning the decision-making policy of an experienced and fixed expert by using its historical behaviors. IL utilizes experts’ demonstrations for policy fitting and aims to extract knowledge from experts’ demonstrations to replicate their behaviors or environments. It can be divided into two main categories: Behavior Cloning (BC) [Bain and Sammut, 1995] and Inverse Reinforcement Learning (IRL) [Russell, 1998]. BC approaches aim to learn a policy that directly maps states to actions, while IRL approaches focus on recovering the reward parameters from demonstrations to recover the experts’ policy [Zheng et al., 2022].
@@
A major challenge in imitation learning (IL) is the substantial requirement of expert demonstration data [Abbeel and Ng, 2004, Hussein et al., 2017]. There is a common assumption that the experts are always experienced (see Figure 1). However, in real-world streaming scenarios, the behavior policy of the expert constantly evolves over time, transitioning from a novice expert to an experienced one. Throughout this process, we accumulate a significant amount of expert’s behavioral evolution history. Taking streaming recommender systems as an example [Zhang et al., 2021, 2022], the agent in streaming recommendation needs to make a trade-off between exploitation and exploration during the recommendation process, and continuously and incrementally update its recommendation policy [Chandramouli et al., 2011, Chang et al., 2017, Jakomin et al., 2020].
Figure 1: Imitation Learning (IL) approaches often hold the view that interactions made by experts are experienced, which may not conform to streaming scenarios since even the expert has to evolve from novice to experienced.
@@
In the behavioral evolution history, there exist a significant amount of contradictory data where an expert may take different actions when facing the same context at different time periods. This contradicts the assumption made by traditional imitation learning (IL) approaches regarding the consistency of expert behavior data. Therefore, directly applying existing IL methods to recover the behavior policy from the expert’s behavioral evolution history is not reasonable. For example, behavior cloning (BC) approaches are highly sensitive to distribution drift and contradictory data, and inconsistencies in the expert’s evolution history can lead to a sharp decline in the final recovery performance. Furthermore, some IL approaches rely on large-scale iterations of Metropolis-Hastings sampling to achieve accurate estimation in universal scenarios [Ramachandran and Amir, 2007, Choi and Kim, 2011]. However, these approaches overlook the exploitation-exploration mechanism in expert interactions and demonstrate low sample efficiency when dealing with large-scale interaction data. While recent work by Hüyük et al. [2022] focused on contextual bandits, it cannot be adapted to the batched bandit setting as it assumes the expert updates the policy in a fully-online manner. Additionally, existing inverse bandit approaches typically employ sampling techniques to estimate the likelihood, resulting in low training efficiency. There is an urgent need to develop inverse bandit methods specifically designed for the more general batched bandit setting and to devise novel techniques for recovering parameters tailored for bandit polices.
@@
In this paper, we focus on the more general batched contextual bandit (BCB) setting, and propose inverse batched contextual bandit (IBCB) to overcome several issues mentioned above. To learn from the evolving behaviors made by online decision-making experts with BCB policies, we design IBCB with a unified framework for both deterministic [Han et al., 2020] and randomized [Dimakopoulou et al., 2019] bandit policies. Under this unified framework, we take linear constraints with expectation relaxation in BCB’s exploitation-exploration pairs and batched updating strategy into consideration. We formulate the inverse problem of IBCB into a simple quadratic programming problem without notifying expert’s behavior rewards. We take the assumption of BCB setting, rather than simply cloning behaviors with no assumption or blindly using large-scale samplings. Combining all these improvements together accelerates the train speed of IBCB and insures the robustness of IBCB for out-of-distribution and contradictory data. Due to that original BCB learns the reward parameter through online learning, IBCB can capture both expert’s policy parameters and reward parameters at the same time.
@@
We summarize our major contributions as follows: (1) We define a new inverse bandit problem with behavioral evolution history in the BCB setting; (2) We introduce a unified framework called IBCB, designed for both deterministic and randomized BCB policies, which can efficiently learn from the interaction history data of novice experts without requiring experts’ feedbacks; (3) IBCB is capable of simultaneously learning the parameters of the expert policy and the reward feedback, which outperforms various existing baseline IL approaches in scenarios such as out-of-distribution data and contradictory data, while also demonstrating significantly improved training speed.
@@ 2 Related Works
Contextual Bandits have been extensively utilized for solving sequential decision-making problems in online learning [Li et al., 2010, Lan and Baraniuk, 2016, Yang et al., 2021]. Recently, there has been increased research focus on a more general setting called batched contextual bandit (BCB) [Han et al., 2020, Ren and Zhou, 2020, Gu et al., 2021], where actions within the same batch share fixed policy parameters. In this regard, Han et al. [2020] proposed a UCB-like update strategy for BCB, which is applicable to both random and adversarial contextual data.
Imitation Learning (IL) is typically used to learn the expert’s action selection policy or estimate the reward parameter from the expert’s behavior [Abbeel and Ng, 2004, Hussein et al., 2017, Arora and Doshi, 2021, Zheng et al., 2022]. IL can be divided into two categories. The first category is Behavior Cloning (BC) [Bain and Sammut, 1995], which directly constructs a direct mapping from state to action [Abbeel et al., 2010, Osa et al., 2017, Liu et al., 2020]. The second category is the inverse reinforcement learning (IRL) method [Russell, 1998], which attempts to recover environment’s reward parameter from the expert’s behavior, that is, imitating the reward parameters. Existing works mainly study IRL based on Bayesian [Ramachandran and Amir, 2007, Choi and Kim, 2011, Balakrishna et al., 2020], maximum entropy (max-ent) [Ziebart et al., 2008, Fu et al., 2018, Qureshi et al., 2019], and generative [Ho and Ermon, 2016, Chen et al., 2021] approaches. Traditional IL approaches mostly assume that the expert’s behavior is optimal at every step [Ziebart et al., 2008, Hussein et al., 2017, Brown and Niekum, 2019], but ignore that early interactions are likely to be the evolutionary data of novice experts learned to become experienced experts.
Offline Reinforcement Learning (Offline RL) has been widely used to improve the model’s ability to generalize to out-of-distribution data, i.e., extrapolation error [Prudencio et al., 2023, Levine et al., 2020]. There are two main research focuses in existing offline RL. One is to restrict the policy during training to avoid out-of-distribution generalization issues, which can be further subdivided into explicitly restricted policies [Mnih et al., 2015, Fujimoto et al., 2019, Kumar et al., 2020, Kostrikov et al., 2021]and implicitly restricted policies [Peng et al., 2019, Chen et al., 2020]. The other is to allow the learning model to adaptively estimate the uncertainty of the environment or data [Kidambi et al., 2020, Yu et al., 2021]. Unlike Offline RL, which requires knowledge of the reward feedback, the proposed IBCB can learn about reward and expert policy’s parameters from the history of expert behavioral evolution without requiring reward feedback information from the environment.
Figure 2: Inverse bandit problem with behavioral evolution history in BCB setting at step $b$ in $n$-th episode, where the rewards are invisible and our goal is to infer and estimate the policy parameters as well as the reward parameters.
@@ 3 Problem Formulation
Batched Contextual Bandit (BCB) Setting: Following the setups in linear contextual bandit literature [Dimakopoulou et al., 2019, Yang et al., 2021, Li et al., 2010], for any context $\bm{s}_{i}\in\mathcal{S}\subseteq\mathbb{R}^{d}$, we assume that the expectation of the observed reward $R_{i}$ from environment is determined by unknown true reward parameters $\bm{\theta}^{*}\in\mathbb{R}^{d}$: $\mathbb{E}[R_{i}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{s}_{i}]=\langle\bm{\theta}^{*},\bm{s}_{i}\rangle.$ In this paper, our focus is on the generalized version of the traditional contextual bandit called the batched contextual bandit (BCB) setting, which has gained considerable attention in the field of bandit theory and applications [Han et al., 2020, Zhang et al., 2021]. For an expert policy in BCB setting, we define expert’s decision-making progress is partitioned into $N$ episodes, and in each episode, expert consists of two phases: (1) the online decision-making chooses the action (i.e., candidate context, and also behavior) for execution from each step’s candidate context set following the updated and fixed expert policy $f$ for $B$ steps ($B$ is also called batch size), and finally stores all context set & action pairs and the observed rewards of the executed actions (contexts) into the data buffer $\mathcal{D}$ (expert’s behavioral evolution history with rewards); (2) the policy updating approximates the optimal policy based on the received actions and rewards in $\mathcal{D}$. Note that for expert policy itself to evolve, expert can observe the reward of executed actions. Also, data buffer stores the indexes of the executed actions rather than actions themselves since each step’s context set contains all actions, including the executed actions. This saves the storage space for practical applications.
Figure 2 shows the content we can observe, where reward of executed actions (behaviors) is unaccessible. Then, we introduce the formal definition of inverse bandit problem with behavioral evolution history in BCB setting as follows.
@@ 4 IBCB: The Proposed Approach
In this section, we proposed Inverse Batched Contextual Bandit approach named IBCB which can efficiently learn from expert’s behavioral evolution history.
Policy Updating in BCB Setting: At the end of the $n$-th episode in the online learning period, the context vectors (corresponding to the executed actions) and their rewards are observed, and are stored into a context matrix $\bm{S}_{n}\in\mathbb{R}^{B\times d}$ and a reward vector $\bm{R}_{n}\in\mathbb{R}^{B}$, respectively. In particular, each row index of $\bm{S}_{n}$ and $\bm{R}_{n}$ corresponds to the index of the step in which the context-reward pair was received. We introduce the updating process of the reward parameter vector $\bm{\theta}_{n}$ in BCB. We first concatenate the context and reward matrices from the previous episodes:
@@
where $\bm{\Phi}_{n+1}=\bm{\Phi}_{n}+\bm{S}_{n}^{\top}\bm{S}_{n}=\sum_{i\in\left[n\right]}\bm{S}_{n}^{\top}\bm{S}_{n},\quad\bm{b}_{n+1}=\bm{b}_{n}+\bm{S}_{n}^{\top}\bm{R}_{n}=\sum_{i\in\left[n\right]}\bm{S}_{n}^{\top}\bm{R}_{n},\quad\bm{\Psi}_{n+1}=\lambda\bm{I}_{d}+\bm{\Phi}_{n+1}.$
@@
2:  $\bm{\Phi}_{n+1}\leftarrow\bm{\Phi}_{n}+\bm{S}_{n}^{\top}\bm{S}_{n},$$\bm{b}_{n+1}\leftarrow\bm{b}_{n}+\bm{S}_{n}^{\top}\bm{R}_{n}$
3:  $\bm{\Psi}_{n+1}\leftarrow\lambda\bm{I}_{d}+\bm{\Phi}_{n+1},$$\bm{\theta}_{n+1}\leftarrow(\bm{\Psi}_{n+1})^{-1}\bm{b}_{n+1}$
@@
5:  $f_{n+1}$ selects the context as $I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\bm{\theta}_{n+1},\bm{s}_{I}\rangle+\alpha[\bm{s}_{I}^{\top}\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{s}_{I}]^{\frac{1}{2}}$, where $\bm{s}_{I}\in\mathcal{S}_{n,b}$
@@
7:  Draw $\tilde{\bm{\theta}}_{n+1}$ from $\mathcal{N}(\bm{\theta}_{n+1},\alpha^{2}(\bm{\Psi}_{n+1})^{-1})$
8:  $f_{n+1}$ selects the context as $I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\tilde{\bm{\theta}}_{n+1},\bm{s}_{I}\rangle$, where $\bm{s}_{I}\in\mathcal{S}_{n,b}$
9:  Return $\bm{\theta}_{n+1},$  $\left(\bm{\Psi}_{n+1}\right)^{-1}$
Finally, as shown in Algorithm 1, the batched UCB policy updates the policy parameters and then selects the context $\bm{s}_{I_{n,b}}$ by applying the policy $f_{n+1}$ according to the following rule: $I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\bm{\theta}_{n+1},\bm{s}_{I}\rangle+\alpha[\bm{s}_{I}^{\top}\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{s}_{I}]^{\frac{1}{2}}$, where $\bm{s}_{I}\in\mathcal{S}_{n,b}$. On the other hand, the batched Thompson sampling policy is a randomized policy that samples the policy parameters from a Gaussian distribution. However, as demonstrated in Theorem 1, we can prove that the batched Thompson sampling policy can also be expressed in the same regularized form as the UCB policy.

### v2

# IBCB: Efficient Inverse Batched Contextual Bandit for Behavioral Evolution History

@@ Abstract
Traditional imitation learning focuses on modeling the behavioral mechanisms of experts, which requires a large amount of interaction history generated by some fixed expert. However, in many streaming applications, such as streaming recommender systems, online decision-makers typically engage in online learning during the decision-making process, meaning that the interaction history generated by online decision-makers includes their behavioral evolution from novice expert to experienced expert. This poses a new challenge for existing imitation learning approaches that can only utilize data from experienced experts. To address this issue, this paper proposes an inverse batched contextual bandit (IBCB) framework that can efficiently perform estimations of environment reward parameters and learned policy based on the expert’s behavioral evolution history. Specifically, IBCB formulates the inverse problem into a simple quadratic programming problem by utilizing the behavioral evolution history of the batched contextual bandit with inaccessible rewards, and it can be extended to fairness-aware expert limitation. We demonstrate that IBCB is a unified framework for both deterministic and randomized bandit policies. The experimental results indicate that IBCB outperforms several existing imitation learning algorithms on synthetic and real-world data and significantly reduces running time. Additionally, empirical analyses reveal that IBCB exhibits better imitation ability for fairness-aware experts, out-of-distribution generalization and is highly effective in learning the bandit policy from the interaction history of novice experts. The code is publicly available.
Index Terms:
batched contextual bandit, imitation learning, behavioral evolution, fairness-aware expert, quadratic programming
@@ 1 Introduction
Traditional imitation learning (IL) focuses on learning the decision-making policy of an experienced and fixed expert by using its historical behaviors. IL utilizes experts’ demonstrations for policy fitting and aims to extract knowledge from experts’ demonstrations to replicate their behaviors or environments. It can be divided into two main categories: Behavior Cloning (BC) [1] and Inverse Reinforcement Learning (IRL) [2]. BC approaches aim to learn a policy that directly maps states to actions, while IRL approaches focus on recovering the reward parameters from demonstrations to recover the experts’ policy [3].
@@
A major challenge in imitation learning (IL) is the substantial requirement of expert demonstration data [4, 5]. There is a common assumption that the experts are always experienced. However, in real-world streaming scenarios, the behavior policy of the expert constantly evolves over time, transitioning from a novice expert to an experienced one. Throughout this process, we accumulate a significant amount of expert’s behavioral evolution history. Taking streaming recommender systems as an example [6, 7], the agent in streaming recommendation needs to make a trade-off between exploitation and exploration during the recommendation process, and continuously and incrementally update its recommendation policy [8, 9, 10].
@@
In the behavioral evolution history, there exist a significant amount of contradictory data where an expert may take different actions when facing the same context at different time periods. This contradicts the assumption made by traditional imitation learning (IL) approaches regarding the consistency of expert behavior data. Meanwhile, recent works like [11, 12] focus on maintaining fairness for bandit algorithms, which introduced a distribution-based random action selection, rather than simply selecting the best action, adding the difficulty to recover expert’s parameters. Therefore, directly applying existing IL methods to recover the behavior policy from the expert’s behavioral evolution history is not reasonable. For example, behavior cloning (BC) approaches are highly sensitive to distribution drift and contradictory data, and inconsistencies in the expert’s evolution history can lead to a sharp decline in the final recovery performance. Furthermore, some IL approaches rely on large-scale iterations of Metropolis-Hastings sampling to achieve accurate estimation in universal scenarios [13, 14]. However, these approaches overlook the exploitation-exploration mechanism in expert interactions and demonstrate low sample efficiency when dealing with large-scale interaction data. While recent work by [15] focused on contextual bandits, it cannot be adapted to the batched bandit setting as it assumes the expert updates the policy in a fully-online manner. Additionally, existing inverse bandit approaches typically employ sampling techniques to estimate the likelihood, resulting in low training efficiency. There is an urgent need to develop inverse bandit methods specifically designed for the more general batched bandit setting and to devise novel techniques for recovering parameters tailored for bandit polices.
@@
In this paper, we focus on the more general batched contextual bandit (BCB) setting, and propose inverse batched contextual bandit (IBCB) to overcome several issues mentioned above. To learn from the evolving behaviors made by online decision-making experts with BCB policies, we design IBCB with a unified framework for both deterministic [16] and randomized [17] bandit policies. Under this unified framework, we take linear constraints with expectation relaxation in BCB’s exploitation-exploration pairs and batched updating policy into consideration. We formulate the inverse problem of IBCB into a simple quadratic programming problem without notifying expert’s behavior rewards. We take the assumption of BCB setting, rather than simply cloning behaviors with no assumption or blindly using large-scale samplings. Combining all these improvements together accelerates the train speed of IBCB and insures the robustness of IBCB for fairness-aware scenarios, out-of-distribution and contradictory data. Due to that original BCB learns the reward parameter through online learning, IBCB can capture both expert’s policy parameters and reward parameters at the same time.
@@
We summarize our major contributions: (1) We define a new inverse bandit problem with behavioral evolution history in the BCB setting; (2) We introduce a unified framework called IBCB, designed for both deterministic and randomized BCB policies, which can efficiently learn from the interaction history data of novice experts without requiring experts’ feedbacks; (3) IBCB is capable of simultaneously learning the parameters of the expert policy and the reward feedback, which outperforms various existing baseline IL approaches in scenarios such as fairness-aware experts, out-of-distribution data and contradictory data, with notable improvements in training speed.
@@ 2 Related Works
Contextual Bandits have been extensively utilized for solving sequential decision-making problems in online learning [18, 19, 20]. Recently, there has been increased research focus on a more general setting called batched contextual bandit (BCB) [16, 21, 22], where actions within the same batch share fixed policy parameters. In this regard, [16] proposed a UCB-like update policy for BCB, which is applicable to both random and adversarial contextual data.
Imitation Learning (IL) is typically used to learn the expert’s action selection policy or estimate the reward parameter from the expert’s behavior [4, 5, 23, 3]. IL can be divided into two categories. The first category is Behavior Cloning (BC) [1], which directly constructs a direct mapping from state to action [24, 25, 26]. The second category is the inverse reinforcement learning (IRL) method [2], which attempts to recover environment’s reward parameter from the expert’s behavior, that is, imitating the reward parameters. Existing works mainly study IRL based on Bayesian [13, 14, 27], maximum entropy (max-ent) [28, 29, 30], and generative [31, 32] approaches. Traditional IL approaches mostly assume that the expert’s behavior is optimal at every step [28, 5, 33], but ignore that early interactions are likely to be the evolutionary data of novice experts learned to become experienced experts.
Offline Reinforcement Learning (Offline RL) has been widely used to improve the model’s ability to generalize to out-of-distribution data, i.e., extrapolation error [34, 35]. There are two main research focuses in existing offline RL. One is to restrict the policy during training to avoid out-of-distribution generalization issues, which can be further subdivided into explicitly restricted policies [36, 37, 38, 39]and implicitly restricted policies [40, 41]. The other is to allow the learning model to adaptively estimate the uncertainty of the environment or data [42, 43]. Unlike Offline RL, which requires knowledge of the reward feedback, the proposed IBCB can learn about reward and expert policy’s parameters from the history of expert behavioral evolution without requiring reward feedback information from the environment.
@@ 3 Problem Formulation
Batched Contextual Bandit (BCB) Setting. Following the setups in linear contextual bandit literature [17, 20, 18], for any context $\bm{s}_{i}\in\mathcal{S}\subseteq\mathbb{R}^{d}$, we assume that the expectation of the observed reward $R_{i}$ from environment is determined by unknown true reward parameters $\bm{\theta}^{*}\in\mathbb{R}^{d}$: $\mathbb{E}[R_{i}~{}|~{}\bm{s}_{i}]=\langle\bm{\theta}^{*},\bm{s}_{i}\rangle.$ In this paper, our focus is on the generalized version of the traditional contextual bandit called the batched contextual bandit (BCB) setting, which has gained considerable attention in the field of bandit theory and applications [16, 6]. For an expert policy in BCB setting, we define expert’s decision-making progress is partitioned into $N$ episodes, and in each episode, expert consists of two phases: (1) the online decision-making chooses the action (i.e., candidate context, and also behavior) for execution from each step’s candidate context set following the updated and fixed expert policy $f$ for $B$ steps ($B$ is also called batch size), and finally stores all context set & action pairs and the observed rewards of the executed actions (contexts) into the data buffer $\mathcal{D}$ (expert’s behavioral evolution history with rewards); (2) the policy updating approximates the optimal policy based on the received actions and rewards in $\mathcal{D}$. Note that for expert policy itself to evolve, expert can observe the reward of executed actions. Also, data buffer stores the indexes of the executed actions rather than actions themselves since each step’s context set contains all actions, including the executed actions. This saves the storage space for practical applications.
Figure 1: Inverse bandit problem with behavioral evolution history in BCB setting at step $b$ in $n$-th episode, where the rewards are inaccessible and our goal is to infer and estimate the policy parameters as well as the reward parameters.
Inverse Bandit Problem. Then, we introduce the formal definition of inverse bandit problem with behavioral evolution history in BCB setting as follows. Figure 1 shows the information that inverse batched contextual bandit (IBCB) can observe, where reward of executed actions (behaviors) is inaccessible. If expert do not update its parameter after each episode, this problem setting will degenerate to the traditional imitation learning problem.
@@ 4 IBCB: The Proposed Approach
In this section, we propose Inverse Batched Contextual Bandit approach named IBCB which can efficiently learn from expert’s behavioral evolution history.
4.1 Policy Updating in BCB Setting
At the end of the $n$-th episode in the online learning period, the context vectors (corresponding to the executed actions) and their rewards are observed, and are stored into a context matrix $\bm{S}_{n}\in\mathbb{R}^{B\times d}$ and a reward vector $\bm{R}_{n}\in\mathbb{R}^{B}$, respectively. In particular, each row index of $\bm{S}_{n}$ and $\bm{R}_{n}$ corresponds to the index of the step in which the context-reward pair was received. We introduce the updating process of the reward parameter vector $\bm{\theta}_{n}$ in BCB. We first concatenate the context and reward matrices from the previous episodes:
@@ 4.1 Policy Updating in BCB Setting
where $\bm{\Phi}_{n+1}=\bm{\Phi}_{n}+\bm{S}_{n}^{\top}\bm{S}_{n},\quad\bm{b}_{n+1}=\bm{b}_{n}+\bm{S}_{n}^{\top}\bm{R}_{n},\quad\bm{\Psi}_{n+1}=\lambda\bm{I}_{d}+\bm{\Phi}_{n+1}.$
@@
2:  $\bm{\Phi}_{n+1}\leftarrow\bm{\Phi}_{n}+\bm{S}_{n}^{\top}\bm{S}_{n},$$\bm{b}_{n+1}\leftarrow\bm{b}_{n}+\bm{S}_{n}^{\top}\bm{R}_{n}$,  $\bm{\Psi}_{n+1}\leftarrow\lambda\bm{I}_{d}+\bm{\Phi}_{n+1},$$\bm{\theta}_{n+1}\leftarrow(\bm{\Psi}_{n+1})^{-1}\bm{b}_{n+1}$
@@
4:  $f_{n+1}$ selects the context as $I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\bm{\theta}_{n+1},\bm{s}_{I}\rangle+\alpha[\bm{s}_{I}^{\top}\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{s}_{I}]^{\frac{1}{2}}$, where $\bm{s}_{I}\in\mathcal{S}_{n,b}$
@@
6:  Draw $\tilde{\bm{\theta}}_{n+1}$ from $\mathcal{N}(\bm{\theta}_{n+1},\alpha^{2}(\bm{\Psi}_{n+1})^{-1})$
7:  $f_{n+1}$ selects the context as $I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\tilde{\bm{\theta}}_{n+1},\bm{s}_{I}\rangle$, where $\bm{s}_{I}\in\mathcal{S}_{n,b}$
8:  Return $\bm{\theta}_{n+1},$  $\left(\bm{\Psi}_{n+1}\right)^{-1}$
Finally, as shown in Algorithm 1, the batched UCB policy updates the policy parameters and then selects the context $\bm{s}_{I_{n,b}}$ by applying the policy $f_{n+1}$ according to the following rule: $I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\bm{\theta}_{n+1},\bm{s}_{I}\rangle+\alpha[\bm{s}_{I}^{\top}\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{s}_{I}]^{\frac{1}{2}}$, where $\bm{s}_{I}\in\mathcal{S}_{n,b}$. On the other hand, the batched Thompson sampling policy is a randomized policy that samples the policy parameters from a Gaussian distribution. However, as demonstrated in Theorem 1, we can prove that the batched Thompson sampling policy can also be expressed in the same regularized form as the UCB policy. Detailed proof can be found in Section 4.1.1.